Paula invited 15 friends to her Sweet 16 birthday party. She decided to use 16 small square tables to form a large square so that she and her friends would all have a place to sit.
At the last minute, four cousins arrived from out of town. How could Paula arrange the 16 tables into a new rectangular shape (with no holes in the middle) so that all 20 people will have a place and there will be no empty seats?
EXTENSION: What is the greatest number of people that can be seated at 16 tables arranged in a rectangular shape?

Answer by stanbon(75887)

Paula invited 15 friends to her Sweet 16 birthday party. She decided to use 16 small square tables to form a large square so that she and her friends would all have a place to sit.
At the last minute, four cousins arrived from out of town. How could Paula arrange the 16 tables into a new rectangular shape (with no holes in the middle) so that all 20 people will have a place and there will be no empty seats?
----------
Draw the pictures of the rectangular arrangements.
A 4 by 4 seats 4+4+4+4 = 16
---
A 2 by 8 seats 2+2+8+8 = 20
---
A 1 by 16 seats 1+1+16+16=34
======================================
EXTENSION: What is the greatest number of people that can be seated at 16 tables arranged in a rectangular shape? 34
============================================
Cheers,
Stan H.